(5 points) Rank the following gases from least dense to most dense at 1.00 atm and 298 K: CO, O2, NO, Ch, HF. Explain (H: 1.00 g/mol; C: 12.0 g/mol; O: 16.0 g/mol; N: 14.0 g/mol; Cl: 35.5 g/mol; 9. F: 19.0 g/mol)
Answer #1

Apply Ideal Gas Law,

PV = nRT

where

P = absolute pressure; V = total volume of gas

n = moles of gas

T = absolute Temperature; R = ideal gas constant

D = mass / volume and Molar volume = n/V

MW = molar weight of sample

Then,

PV = nRT; get molar volume

n/V = P/(RT)

get mass -> n = mass/MW

mass/(MW*V) = P/(RT)

D = mass/V

D/MW = P/(RT)

D = P*MW/(RT)

clearly

at same conditions, this depends only on molar mass

so

CO = 28, O2 = 32, N2O = 44, Cl2 = 70, HF = 20

least density to most dense

HF < CO < O2 < N2O < Cl2